Suppose the line through points (-1,6) and (x,2) is perpendicular to the graph of 2x + y = 3. Find the value of x
Found 2 solutions by nerdybill, solver91311:
Answer by nerdybill(7384)

Suppose the line through points (-1,6) and (x,2) is perpendicular to the graph of 2x + y = 3. Find the value of x
.
Find slope of
2x + y = 3
by rearranging into the "slope-intercept" form:
2x + y = 3
y = -2x + 3
slope then, is -2
.
A line perpendicular to this must have a slope that is the "negative reciprocal":
-2m = -1
m = 1/2
.
slope of (-1,6) and (x,2)
(y2-y1)/(x2-x1) = 1/2
(2-(-6))/(x-(-1)) = 1/2
(2+6)/(x+1) = 1/2
8/(x+1) = 1/2
cross-multiply:
16 = x+1
15 = x

Answer by solver91311(24713)



Step 1: Determine the slope of the line represented by the given equation.

Step 2: Determine the slope of any line perpendicular to the line represented by the given equation using the fact that perpendiculars have slopes that are negative reciprocals:



Step 3: Use the slope formula to determine an expression in for the slope of the line that passes through the two given points.



where and are the coordinates of the given points.

Step 4: Set the expression derived in Step 3 equal to the slope value determined in Step 2. Solve for .

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism